OpenStudy (anonymous):

what is the least common denominator for the numbers 15, 3 ,2 and -6?

jhonyy9 (jhonyy9):

30 because 15*2=30 3*10=30 2*15=30 (-6)*(-5)=30

OpenStudy (anonymous):

Thanks fo the help!

jhonyy9 (jhonyy9):

my pleasure